OpenStudy (savannahlanier40):

Solve the quadratic equation. 3x2 - 14x - 5 = 0 A) x = -3, and x = 5 B) x = 3, and x = -5 C) x = - 1 3 , and x = 5 D) x = 1 3 , and x = -5

OpenStudy (savannahlanier40):

help

OpenStudy (jjcoolja12):

Let's solve your equation step-by-step. 3x2−14x−5=0 Step 1: Factor left side of equation. (3x+1)(x−5)=0 Step 2: Set factors equal to 0. 3x+1=0 or x−5=0 x= −1 3 or x=5 Answer: x= −1 3 or x=5

OpenStudy (jjcoolja12):

http://www.mathpapa.com/algebra-calculator.html

alones (alones):

@jjcoolja12 that would be better if you guide them throu the work, not just show it.Please guy the user thrue the work. Thanks!